7.- -Calculate the flow rate in milliliters per hour for a patient receiving an IV of 250 milliliters of 5% D/W with 250 milligrams of aminophylline. The medication order is 0.006 milligrams per kilogram per minute and the patient weighs 66 kilograms. The physician ordered cisplatin 20 mg/m2. IV for a patient with BSA of 1.7 m2. The label of cisplatin reads 10mg- 10mL 8. How many milligrams should the patient receive? b.How many milliters contain he prescribed dose? Add the correct amount of cisplatin to 1000 mL of 0.22% NS and calculate the rate of flow in milliliters per hour if you infuse in 8 hours. C. The prescriber has ordered 250 milliliters of 5% D/W with 2500 units of heparin. The patient is to receive 50 milliliters per hour IV. 9. How many units per hour will your patient receive? 10. The order states that 200 milligrams of morphine are to be added to 250 milliliters of 5% D/W, and this solution is to be infused at a rate of 10 milligrams per hour. a. How many milliliters per hour will your patient receive? 11. The medication order is for 180 milligrams of morphine sulfate to be added to 250 milliliters of 5% D/V. Your patient is to receive 0.005 milligrams per kilogram per minute IV. How many milligrams per hour should a patient who weighs 100 pounds receive? 12. The prescriber ordered: amikacin sulfate 120 mg/m2/hr IV. The package insert directions are as follows: Add 250 mg to 100 mL 5% D/W Calculate the flow rate in milliliters per hour if the patients BSA is 0.9 square meters.
Answer #1

7.

volume=250 ml

we have total dose of aminophyllin= 250 mg

ordered dose = 0.006 ml / kg/min

weight of patient= 66 kg

total dose for the patient / hour= (0.006*66 kg)*60= 23.76 mg =24 mg/ hr

if 250 mg aminophylline diluted in 250ml solution, 1 ml= 1mg of aminophylline

thus ,flow rate for 1 hour= 24 ml/hr

8

dose ordered= 20 mg/m2

a.

basal surface area= 1.7 m2

so, ordered dose in milligram =20*1.7=34 mg

b. given 10 mg= 10 ml

so ordered dose in millilitre = 34 ml

c. ordered dose = 34 mg

total volume to infuse = 1000ml

rate of flow in 1 hour= 1000/8 =129.25 ml/hr

9.

volume to be infuse= 250 ml

dose to be added= 2500 units

infusion rate per hour= 50 ml

so 1 ml =2500/250= 10 units

thus in 50 ml =50*10= 500 units

10.

dose to be added= 200 milligrams

volume to be infused= 250 ml

infusion rate per hour= 10 milligrams

1ml =200/250= 0.8 mg

millilitre per hour to infuse =10/0.8 =12.5 ml

(1:x=0.8:10

x= 10/0.8 =12.5 ml)